What is the explicit formula for this geometric sequence?

f 1 2 3 4 5 6
f(n) 1 2 4 8 16 32
1
Expert's answer
2019-09-19T09:24:01-0400

So the values we have are:

n : 1 2 3 4 5 6

and

f(n) : 1 2 4 8 16 32


Thus, the explicit formula for this geometric sequence is :

f(n) = "2^{n-1}"


Thus, on substituting the values of n provided to us:


When n = 1, f(1) = "2^{1-1} = 2^0 = 1"

When n = 2, f(2) = "2^{2-1} = 2^1 = 2"

When n = 3, f(3) = "2^{3-1} = 2^2 = 4"

and so on...